Grade 12Magical Mathematics[Interesting Approach]

Find the no of solutions of the equation
2^x = x^2 + 1
Caution : answer is not 2 ;)

Roshan Mohanty
13 Years agoGrade 12

2 Answers

Vikas TU
13 Years ago

2^x = x^2 + 1

Take log both sides with base ''e''

xlog2 = log(x^2 + 1)

x = log(x^2 + 1)/log2

x = log[(x^2 + 1)/(2)]

(x^2 + 1)/(2) = e^x 

x^2/2 + 1/2   = e^x                                  

 

 

 

From this graph we conclude that both graph cuts each other at only one point.

Hence there is onle and only one solution.

 

 

Plz Approve!

 


Roshan Mohanty
13 Years ago

No no no see
u will surely get two values 1 and 0
Now if u take f(x)=x^2 + 1 - 2^x
find f''(x) now 
apply rolles theorem
u will get that f(x)=0 has 3 roots
another root will be b/w 4 and 5